Final Exam 1/13/19 Nam If two distributions have the same range, then what must be a) their high and low data values b) the entire distributions (e) their standard deviations d) the difference between their high and low data values e) their means 1. In a skewed distribution, the median is preferred over the mean as a measure of the center because a) the median is the most frequent number b) the mean may be too heavily influenced by outliers 2. the median is less than the mean e) the median measures the arithmetic average of the data excluding outliers d) the mean measures the spread in the data If a distribution has a mean of 9 and five of the six data values are 6, 8, 11, 15, and 18, then the missing data value is: a) 0 c)-4 d)-6 e) none of these Gail received a grade of 75 on her chemistry exam. Rich received a grade of 80 on his law exam. The mean for the chemistry exam is 70 with a standard deviation of 5 and the mean for the law exam is 75 with a standard deviation of 4. Whose a) Richs, since his grade of 80 is higher than Gails 75 b) Gails, since the class mean in Chemistry is lower than the class mean in law 4. grade is more outstanding? Richs, since he has a higher z score d) Gails, since she has a higher z score e) none of these 53 9 65 71 83 95 101 1.0-5 0 1.0 2.0 2.5 40 50 75 99 raw score percentile rank Qi) 5. Whats the missing raw score? 6. Whats the missing z-score? 7. Whats the difference between the two missing percentile ranks? °) 71 b) 1.0 c 5 b) 5 a)70 c)77 d)95 e)98 1.5 e) 5 c) 10 d) 15 8. Since the mean is to the a) left, positively b) left, negatively (right, positively d) right, negatively e) none of these In a symmetric distribution, what is the approximate percentage of data values that lie between two of the median, this distribution is skewed 9. and one standard deviation above the mean? 81.5% d) 68% e) 34% standard deviations below th a)84% b) 95% Using this set of data: (122, 134, 126, 120, 128, 130, 120, 118, 125, 122, 126, 136, 118, 122, 124, 119) percentile rank of 126 rounded to the nearest c) 78 what is the none of these d) 8I Find the probability of getting a sum - (S or 6) when rolling e pair of dice. a) 5/18 11. b) 5/12 c) 1/4
Answer #1

1)

Range = Highest - Lowest

option D)

2)

option B) mean maybe heavily influenced by outliers

3)

missing value = 9 *6 - (6 + 8 +11 + 15 +18)

=   -4

option C) is correct

4)

z = (X - mean)/sd

z-score for Gail = (75 - 70)/5 = 1

z-score for Rich = (80 - 75)/4 = 1.25

since

Rich > Gail

option C) is correct

5)

X = mean + z* sd

83 = mean + 1 * sd

95 = mean + 2 * sd

solving mean = 71

sd = 12

z = 0

X = mean = 71

option B)

6)

z = (X - mean)/sd = (53 - 71)/12

=   -1.5

7)

percentile = 100 * P(-1.5 < Z < -0.5)

= 25 (approx)